the width and length of a rectangle are consecutive odd integers. If the width is w, find the perimeter, in simplest form, in terms of w.
Answer by lwsshak3(11628)

the width and length of a rectangle are consecutive odd integers. If the width is w, find the perimeter, in simplest form, in terms of w.
**
given width=W
length, L=W+2
perimeter=2(L+W)
=2(W+2+W)
=2(2+2W)
perimeter=4+4W